OpenStudy (anonymous):

What are the foci of (x+10)^2/25-(y+1)^2/9=1? (-1 ±√(34) , -10) (-1, -10 ±√(34)) (-10 ±√(34) , -1) (-10, -1 ±√(34))

OpenStudy (agent0smith):

This time it's a horizontal hyperbola, with center (h, k) which is (-10, -1). Foci of a horizontal hyperbola: \[(h \pm c, k)\] c is the distance from the center to the focus, and is given by c^2 = a^2 + b^2

OpenStudy (agent0smith):

a^2 = 25, b^2 = 9

OpenStudy (agent0smith):

Standard form for horizontal hyperbola is: \[\frac{ (x-h)^2 }{ a^2 } - \frac{ (y-k)^2 }{ b^2 } = 1\]